With given vectors u=<3,-1,-2> v=<-1,0,2> w=<-6,1,4> compute the expressions below
2.1 v+3w
2.2 u-2v
2.3 - (v+3u)
1
Expert's answer
2021-06-16T04:57:11-0400

2.1


"v+3w=\langle-1,0,2\rangle+3\langle-6,1,4\rangle"

"=\langle-1+3(-6), 0+3(1), 2+3(4)\rangle"




"=\langle-19,3, 14\rangle"

2.2


"u-2v=\langle3,-1,-2\rangle+(-2)\langle-1,0,2\rangle"

"=\langle3+(-2)(-1), -1+(-2)(0), -2+(-2)(2)\rangle"

"=\langle5,-1, -6\rangle"



2.3


"-(v+3u)=-(\langle -1,0,2\rangle+3\langle 3,-1,-2\rangle)"

"=-(\langle-1+3(3), 0+3(-1), 2+3(-2)\rangle)"

"=\langle-8,3, 4\rangle"